OpenStudy (anonymous):

Show whether Rolle's Theorem can be applied to f(x) If it can, determine any values of c in the interval [0,2pi] for which f'(c)=0 f(x)=cos x

ganeshie8 (ganeshie8):

wats the hypothesis of Rolle's theorem ?

ganeshie8 (ganeshie8):

we need to show/state f(x) = cosx satisfies the hypothesis of Rolle's theorem right ?

OpenStudy (anonymous):

yes, to show it can be applied. hold on. I'll post it up

ganeshie8 (ganeshie8):

take ur time :)

OpenStudy (anonymous):

Let f be continuous on the closed interval [a,b] and be differentiable on the open interval (a,b). If: f(a)=f(b), then there is at least one number c in (a,b) such that f'(c)=0

OpenStudy (anonymous):

thats it :)

ganeshie8 (ganeshie8):

Yes, lets see if f(x) satisfies the conditions required for applying Rolle's theorem :- ``` Let f be continuous on the closed interval [a,b] and be differentiable on the open interval (a,b). If: f(a)=f(b), ```

ganeshie8 (ganeshie8):

\(f(x) =\cos x\) \(f(x)\) is differentiable in \((0, 2\pi)\) Also, \(f(0) = f(2\pi) = 1\) So, the hypothesis of Rollle's theorem are satisfied.

ganeshie8 (ganeshie8):

next, find f'(x) and set it equal to 0 and solve

ganeshie8 (ganeshie8):

f'(x) = -sin(x) = 0 x = ?

OpenStudy (anonymous):

positive sin?

OpenStudy (anonymous):

x=sin ?

ganeshie8 (ganeshie8):

f'(x) = -sin(x) = 0 x = arcsin(0) = 0, pi, 2pi

ganeshie8 (ganeshie8):

so the values of c for which f'(c) = 0 in interval [0,2pi] are : c = 0, pi, 2pi

OpenStudy (anonymous):

oh right. forgot thats how you did it

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

ganeshie you are the best

OpenStudy (anonymous):

thank you so much

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

just a quick question :- Is the given interval in question really closed ? [0,2pi]

ganeshie8 (ganeshie8):

usually, in questions involving Rolle's theorem, they give open interval (0, 2pi)

OpenStudy (anonymous):

yeah it is. why?

OpenStudy (anonymous):

no it has the brackets

ganeshie8 (ganeshie8):

if it is open interval the answer changes, thats why asking...

OpenStudy (anonymous):

[ ] those little guys

OpenStudy (anonymous):

ah I see

ganeshie8 (ganeshie8):

if it is open interval, answer wud be just : c = pi okay cool :) just want u see thats all

OpenStudy (anonymous):

got it. Thanks. I have another question can you help me with it?

ganeshie8 (ganeshie8):

sure... il try post it :)